# references.bib

### v1

@@ 5.4 Data Set Size Requirements
To evaluate how much data is required to train TeVAE to a point of adequate anomaly detection performance, it has been trained with $1$h, $8$h, $64$h, and $512$h of dynamic testing time. The results for this experiment are presented in Table 4.
@@
$1$h $0.20\pm 0.02$ $0.11\pm 0.01$ $\textbf{0.83}\pm 0.06$ $0.57\pm 0.02$ $321.1\pm 30.4$ $0.07\pm 0.01$
$8$h $0.37\pm 0.14$ $0.26\pm 0.13$ $0.80\pm 0.06$ $0.69\pm 0.04$ $\textbf{298.7}\pm 29.2$ $0.16\pm 0.08$
$64$h $0.69\pm 0.06$ $0.76\pm 0.15$ $0.66\pm 0.03$ $\textbf{0.71}\pm 0.01$ $371.3\pm 38.4$ $0.61\pm 0.15$
$512$h $\textbf{0.70}\pm 0.03$ $\textbf{0.92}\pm 0.08$ $0.57\pm 0.04$ $0.66\pm 0.04$ $418.1\pm 17.3$ $\textbf{0.63}\pm 0.18$
$1$h $0.61\pm 0.05$ $0.72\pm 0.04$ $0.55\pm 0.09$ $0.57\pm 0.02$ $459.9\pm 42.5$ $0.54\pm 0.07$
$8$h $0.74\pm 0.04$ $0.94\pm 0.02$ $0.61\pm 0.06$ $0.69\pm 0.04$ $460.9\pm 35.1$ $0.68\pm 0.04$
$64$h $\textbf{0.77}\pm 0.02$ $\textbf{1.00}\pm 0.00$ $\textbf{0.62}\pm 0.02$ $\textbf{0.71}\pm 0.01$ $418.8\pm\phantom{0}4.5$ $\textbf{0.81}\pm 0.05$
$512$h $0.72\pm 0.03$ $0.97\pm 0.04$ $0.58\pm 0.06$ $0.66\pm 0.04$ $\textbf{412.9}\pm 25.0$ $0.67\pm 0.15$
@@
On the one hand, as the training and validation subset increases in size, the precision value improves but on the other hand recall value decreases as the subset grows, though at a smaller scale compared to the increase in precision. This can be attributed to the fact that smaller subset sizes lead to a small validation set and therefore less data to obtain a threshold from. With a limited amount of data the validation set distribution is very different to the true data distribution, leading to a threshold that is very small and hence marks most anomalies correctly but also leads to a lot of false positives. Despite the decreasing recall, the $F_{1}$ score increases with a growing training and validation subset size. It can also be observed that both the $F_{1,\text{best}}$ and the $A_{\text{PR}}$ reach a point of diminishing returns after $8$h of dynamic testing. Given that neither metric relies on the unsupervised threshold, it indicates that it plays a large role in the $F_{1}$ score. It further implies that the model quality remains largely the same from $8$h onwards. Also, the $F_{1}$ score seems to approach the $F_{1,\text{best}}$ score as the subset grows, also backing the fact that with a small subset size, a good threshold cannot easily be obtained. Therefore, for application on the test benches in automotive powertrain development, the largest subset size is desirable due to the higher precision value and a closer-to-ideal threshold value. When using the unsupervised threshold, the average detection delay increases with a larger training and validation subset and appears to have a relationship with the recall figure. This is due to the fact that $N_{\text{tp}}$ increases and $N_{\text{fn}}$ decreases accordingly, which means that the maximum delay $\delta=T$ is applied more often, leading to the larger average detection delay. Lastly, the root-cause precision $P_{\text{rc}}$ also appears to have a positive correlation with the precision figure. This is to be expected since a root-cause true positive can only occur when an anomaly is correctly identified. However, there is still clearly some room for improvement as $P_{\text{rc}}$ is always lower than or equal to $P$, where in case of a perfect root-cause analysis $P_{\text{rc}}=P$.
@@ 5.6 Hyperparameter Optimisation
$1$ $64$ $\textbf{0.70}\pm 0.03$ $0.92\pm 0.08$ $\textbf{0.57}\pm 0.04$ $\textbf{0.66}\pm 0.04$ $\textbf{418.1}\pm 17.3$ $0.63\pm 0.18$
$8$ $64$ $0.69\pm 0.01$ $0.98\pm 0.03$ $0.54\pm 0.02$ $0.65\pm 0.02$ $446.7\pm 36.9$ $0.66\pm 0.06$
@@
$1$ $64$ $\textbf{0.72}\pm 0.03$ $0.97\pm 0.04$ $\textbf{0.58}\pm 0.06$ $\textbf{0.66}\pm 0.04$ $412.9\pm 25.0$ $0.67\pm 0.15$
$8$ $64$ $0.71\pm 0.01$ $0.92\pm 0.08$ $\textbf{0.58}\pm 0.04$ $0.65\pm 0.02$ $\textbf{411.7}\pm\phantom{0}5.3$ $0.62\pm 0.02$
@@
$1$ $1$ $0.34\pm 0.18$ $\textbf{1.00}\pm 0.00$ $0.22\pm 0.15$ $0.42\pm 0.11$ $690.5\pm 106.8$ $0.74\pm 0.15$
$1$ $8$ $\textbf{0.75}\pm 0.03$ $0.95\pm 0.04$ $\textbf{0.62}\pm 0.06$ $\textbf{0.71}\pm 0.05$ $\textbf{407.8}\pm 24.1$ $0.55\pm 0.10$
$1$ $64$ $0.70\pm 0.03$ $0.92\pm 0.08$ $0.57\pm 0.04$ $0.66\pm 0.04$ $418.1\pm 17.3$ $0.63\pm 0.18$
$1$ $512$ $0.70\pm 0.05$ $0.98\pm 0.03$ $0.55\pm 0.05$ $0.64\pm 0.03$ $448.9\pm 38.0$ $\textbf{0.67}\pm 0.08$
$1$ $1$ $0.49\pm 0.10$ $0.74\pm 0.29$ $0.44\pm 0.12$ $0.42\pm 0.11$ $655.3\pm 99.8$ $0.58\pm 0.27$
$1$ $8$ $\textbf{0.77}\pm 0.03$ $0.94\pm 0.05$ $\textbf{0.65}\pm 0.06$ $\textbf{0.71}\pm 0.05$ $\textbf{387.6}\pm 37.6$ $0.48\pm 0.08$
$1$ $64$ $0.72\pm 0.03$ $0.97\pm 0.04$ $0.58\pm 0.06$ $0.66\pm 0.04$ $412.9\pm 25.0$ $\textbf{0.67}\pm 0.15$
$1$ $512$ $0.72\pm 0.03$ $\textbf{1.00}\pm 0.00$ $0.56\pm 0.04$ $0.64\pm 0.03$ $439.7\pm 24.2$ $0.66\pm 0.04$
@@ 5.7 Benchmarking
Of course, TeVAE is not the first model proposed for time-series anomaly detection. To relate its anomaly detection performance, it is compared with a series of other models based on variational autoencoders. The chosen subset of models is based on the work discussed in Section 3 which either linked source code or contained enough information for implementation. The models are implemented using hyperparameters specified in their respective publications. All models are trained on the $512$h subset with early stopping, which is parameterised equally across all models. The anomaly detection process specified in Algorithm 1 is also applied to all models, along with the threshold estimation method. For VASP and LW-VAE no root-cause precision $P_{\text{rc}}$ is provided because the resulting anomaly scores cannot be decomposed. The results can be seen in Table 8.

### v2

@@ 5.4 Data Set Size Requirements
To evaluate how much data is required to train TeVAE to a point of adequate anomaly detection performance, it has been trained with 11h, 88h, 6464h, and 512512h of dynamic testing time. The results for this experiment are presented in Table 4.
@@
11h $0.20\pm 0.02$ $0.11\pm 0.01$ $\textbf{0.83}\pm 0.06$ $0.57\pm 0.02$ $321.1\pm 30.4$ $0.07\pm 0.01$
88h $0.37\pm 0.14$ $0.26\pm 0.13$ $0.80\pm 0.06$ $0.69\pm 0.04$ $\textbf{298.7}\pm 29.2$ $0.16\pm 0.08$
6464h $0.69\pm 0.06$ $0.76\pm 0.15$ $0.66\pm 0.03$ $\textbf{0.71}\pm 0.01$ $371.3\pm 38.4$ $0.61\pm 0.15$
512512h $\textbf{0.70}\pm 0.03$ $\textbf{0.92}\pm 0.08$ $0.57\pm 0.04$ $0.66\pm 0.04$ $418.1\pm 17.3$ $\textbf{0.63}\pm 0.18$
11h $0.61\pm 0.05$ $0.72\pm 0.04$ $0.55\pm 0.09$ $0.57\pm 0.02$ $459.9\pm 42.5$ $0.54\pm 0.07$
88h $0.74\pm 0.04$ $0.94\pm 0.02$ $0.61\pm 0.06$ $0.69\pm 0.04$ $460.9\pm 35.1$ $0.68\pm 0.04$
6464h $\textbf{0.77}\pm 0.02$ $\textbf{1.00}\pm 0.00$ $\textbf{0.62}\pm 0.02$ $\textbf{0.71}\pm 0.01$ $418.8\pm\phantom{0}4.5$ $\textbf{0.81}\pm 0.05$
512512h $0.72\pm 0.03$ $0.97\pm 0.04$ $0.58\pm 0.06$ $0.66\pm 0.04$ $\textbf{412.9}\pm 25.0$ $0.67\pm 0.15$
@@
On the one hand, as the training and validation subset increases in size, the precision value improves but on the other hand recall value decreases as the subset grows, though at a smaller scale compared to the increase in precision. This can be attributed to the fact that smaller subset sizes lead to a small validation set and therefore less data to obtain a threshold from. With a limited amount of data the validation set distribution is very different to the true data distribution, leading to a threshold that is very small and hence marks most anomalies correctly but also leads to a lot of false positives. Despite the decreasing recall, the $F_{1}$ score increases with a growing training and validation subset size. It can also be observed that both the $F_{1,\text{best}}$ and the $A_{\text{PR}}$ reach a point of diminishing returns after 88h of dynamic testing. Given that neither metric relies on the unsupervised threshold, it indicates that it plays a large role in the $F_{1}$ score. It further implies that the model quality remains largely the same from 88h onwards. Also, the $F_{1}$ score seems to approach the $F_{1,\text{best}}$ score as the subset grows, also backing the fact that with a small subset size, a good threshold cannot easily be obtained. Therefore, for application on the test benches in automotive powertrain development, the largest subset size is desirable due to the higher precision value and a closer-to-ideal threshold value. When using the unsupervised threshold, the average detection delay increases with a larger training and validation subset and appears to have a relationship with the recall figure. This is due to the fact that $N_{\text{tp}}$ increases and $N_{\text{fn}}$ decreases accordingly, which means that the maximum delay $\delta=T$ is applied more often, leading to the larger average detection delay. Lastly, the root-cause precision $P_{\text{rc}}$ also appears to have a positive correlation with the precision figure. This is to be expected since a root-cause true positive can only occur when an anomaly is correctly identified. However, there is still clearly some room for improvement as $P_{\text{rc}}$ is always lower than or equal to $P$, where in case of a perfect root-cause analysis $P_{\text{rc}}=P$.
@@ 5.6 Hyperparameter Optimisation
11 $64$ $\textbf{0.70}\pm 0.03$ $0.92\pm 0.08$ $\textbf{0.57}\pm 0.04$ $\textbf{0.66}\pm 0.04$ $\textbf{418.1}\pm 17.3$ $0.63\pm 0.18$
88 $64$ $0.69\pm 0.01$ $0.98\pm 0.03$ $0.54\pm 0.02$ $0.65\pm 0.02$ $446.7\pm 36.9$ $0.66\pm 0.06$
@@
11 $64$ $\textbf{0.72}\pm 0.03$ $0.97\pm 0.04$ $\textbf{0.58}\pm 0.06$ $\textbf{0.66}\pm 0.04$ $412.9\pm 25.0$ $0.67\pm 0.15$
88 $64$ $0.71\pm 0.01$ $0.92\pm 0.08$ $\textbf{0.58}\pm 0.04$ $0.65\pm 0.02$ $\textbf{411.7}\pm\phantom{0}5.3$ $0.62\pm 0.02$
@@
11 11 $0.34\pm 0.18$ $\textbf{1.00}\pm 0.00$ $0.22\pm 0.15$ $0.42\pm 0.11$ $690.5\pm 106.8$ $0.74\pm 0.15$
11 88 $\textbf{0.75}\pm 0.03$ $0.95\pm 0.04$ $\textbf{0.62}\pm 0.06$ $\textbf{0.71}\pm 0.05$ $\textbf{407.8}\pm 24.1$ $0.55\pm 0.10$
11 $64$ $0.70\pm 0.03$ $0.92\pm 0.08$ $0.57\pm 0.04$ $0.66\pm 0.04$ $418.1\pm 17.3$ $0.63\pm 0.18$
11 $512$ $0.70\pm 0.05$ $0.98\pm 0.03$ $0.55\pm 0.05$ $0.64\pm 0.03$ $448.9\pm 38.0$ $\textbf{0.67}\pm 0.08$
11 11 $0.49\pm 0.10$ $0.74\pm 0.29$ $0.44\pm 0.12$ $0.42\pm 0.11$ $655.3\pm 99.8$ $0.58\pm 0.27$
11 88 $\textbf{0.77}\pm 0.03$ $0.94\pm 0.05$ $\textbf{0.65}\pm 0.06$ $\textbf{0.71}\pm 0.05$ $\textbf{387.6}\pm 37.6$ $0.48\pm 0.08$
11 $64$ $0.72\pm 0.03$ $0.97\pm 0.04$ $0.58\pm 0.06$ $0.66\pm 0.04$ $412.9\pm 25.0$ $\textbf{0.67}\pm 0.15$
11 $512$ $0.72\pm 0.03$ $\textbf{1.00}\pm 0.00$ $0.56\pm 0.04$ $0.64\pm 0.03$ $439.7\pm 24.2$ $0.66\pm 0.04$
@@ 5.7 Benchmarking
Of course, TeVAE is not the first model proposed for time-series anomaly detection. To relate its anomaly detection performance, it is compared with a series of other models based on variational autoencoders. The chosen subset of models is based on the work discussed in Section 3 which either linked source code or contained enough information for implementation. The models are implemented using hyperparameters specified in their respective publications. All models are trained on the 512512h subset with early stopping, which is parameterised equally across all models. The anomaly detection process specified in Algorithm 1 is also applied to all models, along with the threshold estimation method. For VASP and LW-VAE no root-cause precision $P_{\text{rc}}$ is provided because the resulting anomaly scores cannot be decomposed. The results can be seen in Table 8.